A transportation company plans to buy tires in volume. They are choosing between two brands of tires, A and B. In order to come up with a wise decision, a survey was conducted between these two brands of tires.

$$\begin{matrix} & \text{BRAND A} & \text{BRAND B}\\ \text{average mileage} & 40,000\ km & 38,000\ km\\ \text{standard deviation} & 4,500\ km & 3,900\ km \end{matrix} $$

Is Brand A better than Brand B; using a 0.05 level?


1
Expert's answer
2021-02-24T07:40:41-0500

Let us test the null hypothesis "H_0:\mu_1=\mu_2" against the alternative hypothesis "H_1:\mu_1>\mu_2" .

We use the statistic "Z=\frac{\bar{X}_A-\bar{X}_B}{\sqrt{\sigma_x^2+\sigma_y^2}}". If "X_A" and "X_B" are normally distributed, then Z has (approximately) normalized Gauss distribution, and the Laplace function "\Phi(x)" is its CDF (cumulative distribution function). The critical domain corresponding to a statistical significance "\alpha=0.05" is "\Phi(Z)>1-\alpha=0.95" or "Z>\Phi^{-1}(0.95)=1.64"

"Z=\frac{40-38}{\sqrt{4.5^2+3.9^2}}=0.336<1.64"

Therefore, the hypothesis that "\mu_1=\mu_2" can not be rejected at the significance level 0.05.